A large conducting plane has a surface charge density 1*10^-4 C/m2. Find the electrostatic energy stored in a cubical volume of edge 1 cm in front of the plane.

Grade:

1 Answers

Amitabh Shrivastava
13 years ago

Energy=1/2(Epsilon)*E^2

E=1/2(Sigma)(Epsilon)

Energy=1/8*(Epsilon)^3*(Sigma)^2